Find three consecutive even integers such that three times the first equals the sum of the other two
Answer by math_helper(2461)

Let n be the smallest of the 3 consecutive even integers.
3n = (n+2) + (n+4)
3n = 2n + 6
n = 6

Ans: The 3 integers are 6, 8, and 10

Check:
"three times first" —> 3(6) = 18
"sum of other two" —> 8+10 = 18 (ok)